Circular No. 5939

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

NOVA CASSIOPEIAE 1993 S. Boyle, University of London Observatory, reports: “Nova Cas 1993

was observed with the cassegrain spectrograph on the 0.6-m Allen reflector at Mill Hill during January and February. The most recent spectrum (range 505–675 nm, resolution 1.0 nm), obtained on Feb. 20.82 UT, suggests that the nova is entering (or has entered) its nebular phase. The iron lines seen in earlier spectra were not detected. Emission lines at 557.7 nm, 630.0 nm and 636.4 nm ([O I] forbidden lines) were detected. These lines were observed to be narrower, and stronger relative to Hα, than in a spectrum obtained on Feb. 7. The ‘main’ emission-line due to Hα also narrowed between Feb. 7 and 20. A further emission line, with intensity 0.3 of the main Hα emission line, appears at 659.9 nm in the Feb. 20 spectrum. If this line also arises from Hα, it represents a doppler velocity of ∼ 1700 km/s.”

Further visual magnitude estimates: Feb. 18.21 UT, 10.8 (O. Midtskogen, Tranby, Norway); 18.75, 11.2 (B. H. Granslo, Fjellhamar, Norway); 19.22, 11.4 (P. Schmeer, Bischmisheim, Germany); 19.88, 11.6 (G. Comello, Groningen, Netherlands); 20.06, 12.1 (Granslo); 20.80, 12.3 (G. Poyner, Birmingham, England); 21.10, 12.7 (R. King, Duluth, MN); 22.13, [13.3 (R. Royer, Lakewood, CA); 22.42, 12.7 (P. Dombrowski, Glastonbury, CT); 22.76, 13.9 (Midtskogen); 22.86, 13.5 (S. Meister, Bulach, Switzerland).

1993 FW Continuation to the ephemeris on IAUC 5856, from the orbital ele-

ments on MPEC 1994-D09:

1994 TT α_{2000} δ_{2000} Δ r ϵ β V
Feb. 17 12^{h}35.^{m}03 − 3^{o}40.^{′}4 41.393 42.141 138.^{o}7 0.^{o}9 23.4
27 12 34.47 − 3 36.4 41.289 42.140 148.9 0.7 23.3
Mar. 9 12 33.82 − 3 32.0 41.211 42.140 159.1 0.5 23.3
19 12 33.12 − 3 27.1 41.162 42.140 169.3 0.3 23.3
29 12 32.39 − 3 22.1 41.142 42.140 179.4 0.0 23.2
Apr. 8 12 31.66 − 3 17.0 41.152 42.140 170.6 0.2 23.3
18 12 30.95 − 3 12.2 41.192 42.140 160.6 0.5 23.3
28 12 30.29 − 3 07.6 41.259 42.140 150.6 0.7 23.3
May 8 12 29.71 − 3 03.6 41.353 42.139 140.8 0.9 23.3
18 12 29.22 − 3 00.2 41.469 42.139 131.0 1.0 23.4
28 12 28.85 − 2 57.5 41.604 42.139 121.3 1.2 23.4
June 7 12 28.60 − 2 55.6 41.755 42.139 111.6 1.3 23.4

1994 February 23 Brian G. Marsden